OpenStudy (anonymous):

(Stoichiometry) The annual production of sulfur dioxide from burning coal and fossil fuels, for auto exhaust, and from other activities is about 26 million tons. How much sulfur, present in the original materials, would result in that quantity of SO2? S+O2 --> SO2

OpenStudy (jfraser):

Turn those tons into grams and turn those grams into moles

OpenStudy (jfraser):

\[26*10{^6} tons * (\frac{2000lb}{1ton})*(\frac{454g}{1lb})*(\frac{1mol SO{_2}}{64g SO{_2}})\]